The average fruit fly will lay 384 eggs into rotting fruit. A biologist wants to see if the average will change for flies that have a certain gene modified. The data below shows the number of eggs that were laid into rotting fruit by several fruit flies that had this gene modified. Assume that the distribution of the population is normal.

372, 371, 395, 408, 371, 371, 412, 383, 383, 408, 374, 373, 385

What can be concluded at the the α α = 0.05 level of significance level of significance?

a.)

  1. The null and alternative hypotheses would be:

H0: μ =

H1: μ ≠

b.) The test statistic t =    (please show your answer to 3 decimal places.)

c.) The p-value = (Please show your answer to 4 decimal places.)
Answer #1

Answer)

A)

Ho : u = 384

Ha : u is not equal to 384

B)

test statistics = (sample mean - claimed mean)/(s.d/√n)

So, first we need to find the sample mean and s.d for the given data

Mean = 385.0769

S.d = 15.5909

N = 13

t = (385.0769 - 384)/(15.5909/√13)

t = 0.249

C)

Degrees of freedom is = n-1 = 12

For df 12 and 0.249 test statistics, p-value from t distribution is = 0.8076